Raggs Limited, a clothing firm, determines that the revenue, in rand, of selling x dresses is given by the function R(x) = x(450 βˆ’ 0,050x). It also determines that the total cost, in rand, of producing x dresses is given by the function C(x) = 40 000 + 0,025x2. The marginal profit when x dresses are produced and sold, is given by the function


1
Expert's answer
2022-03-25T05:07:26-0400

"P(x)=R(x)-C(x)=x(450-0.05x)-40000-0.025x^2=-0.075x^2+450x-40000."

"MP=\frac{dP}{dx}=-0.15x+450."